# Explainable Malware Analysis: Concepts, Approaches and Challenges

### v1

@@ IV-C Hardware-Based Malware Approaches
Many researchers have designed hardware-based malware detectors with the assumption that solutions such as anti-virus software can be fooled easily by malicious code. There are many works on hardware-assisted malware detection, but there are few works that are explainable and can provide reasonable interpretations for the predictions in a meaningful manner.
Although there are many studies on hardware-assisted malware detection, there is a distinct lack of research focused on explainability. This highlights a critical need for innovation in the field that goes beyond detection accuracy, aiming for systems that can articulate the rationale behind their detections. In table V, the literature on hardware malware classification and detection using XAI is discussed.
Sheldon [105] conducts a study analyzing hardware traces for malware detection using explainable ML approaches. Figure 11 illustrates the process of generating hardware traces, where both malware-infected and benign files are executed on a cutting-edge state-of-the-art hardware platform, leading to the production of hardware traces. Hardware traces comprise the data stored in caches and registers during the execution of a program. These traces are then processed by an ML model to identify malware presence. Subsequently, the model’s accuracy is evaluated against that of leading-edge ML models. The feedback obtained from this comparison is utilized to refine the accuracy of malware detection.
Figure 11: Hardware Trace Data Generation.
To explore the comprehensive study of hardware trace analysis and the development of an explainable, hardware-assisted malware detection framework, the research by Pan et al. [88] introduces a hardware framework depicted in Figure 12. This framework is structured around three principal activities.
@@
After determining the linear regression coefficients, the focus shifts to interpreting the outcomes, with particular emphasis on identifying the most influential features. Features associated with larger coefficients are flagged as potentially malicious. To benchmark the effectiveness of their method, the authors reference the PREEMPT malware detector for comparison [16]. PREEMPT employs two algorithmic models, i.e., Random Forest (RF) and Decision Tree (DT), utilizing hardware performance counters (HPC) to generate its dataset. However, it does not extensively investigate the dataset’s characteristics, focusing primarily on the interpretation of outcomes. Given the scarcity of research in this area and the need to address malware detection challenges, it is imperative for researchers to engage deeply with the realm of explainable hardware malware detection.
Building upon their previous efforts, the authors [89] harness hardware performance counters and embedded trace buffers to identify the exact locations of malicious activities within a system. They develop DT and RNNs to perform a trade-off between accuracy and efficiency in their detection methodology. Their evaluation, conducted on a broad spectrum of real-world malware datasets, elucidates the interpretability of the RNN model, leveraging linear regression and DT through tree parsing techniques.
@@ IV-D PDF-Based Malware Approach
A Portable Document Format (PDF) file contains text, images, digital signatures, and other elements. Its structure includes a Header, Body, Cross-reference Table, and Trailer, as illustrated in Figure 13. The header of a PDF file is the top section that indicates the version number and file format. The body of the PDF stores all the pertinent data, and it contains a range of objects, including data, text, images, and dictionaries. The cross-reference table in a PDF file includes links to all elements within the document, which facilitates navigation and access. The trailer, which links to the cross-reference table, also contains the EOF marker.
Figure 13: Structure of PDF file.
@@
To classify PDF malware, another research [45] introduces an explainable method named LEMNA, which provides high-fidelity explanations for malware detection. They utilize deep learning models and assess their interpretability using LEMNA. The study also explores feature augmentation, along with synthetic and feature deduction tests. They note that due to sparse input feature vectors affecting local decision boundaries, LIME and other advanced explainable techniques were as ineffective as traditional feature selection methods. The PDF-based malware detection studies are summarized in Table V.
@@ IV-E Linux-Based Malware Approach
Wang et al. [119] focus on exposing vulnerabilities in malware detectors through explainability-guided evasion attacks that combine feature space manipulation with problem space obfuscation. They utilize a dataset of approximately 43,553 ELF binary files on Linux systems. Their research uses the model-agnostic explainability method SHAP to demonstrate how evasion attacks can be transferred from one detector to another. In another study, Mills et al. [83] develop a lightweight malware detection system named NODENS, suitable for deployment on Raspberry Pi hardware. They test several ML-based algorithms on a Linux operating system, with the Random Forest algorithm performing optimally among them. This work utilizes a tree-based model to facilitate visual interpretation of the classification process, which enhances the end user’s understanding of the output and aids in the individual development of the malware sample lifecycle. Due to the infrequent application of the Linux platform, there is limited research on explainable malware detection within the Linux domain. Studies addressing this topic are detailed in Table V.
@@ IV-F Other Approaches
TABLE VI: Research Addressing Explainable Machine Learning in Malware Analysis.
@@
Chen 2018  [23] Aims to provide an image-based malware approach for explainability in static malware classification The grayscale images were interpreted and the most relevant attributes that contributed to the training model were shown. Visual interpretations are shown to get interpretation results It has not been demonstrated that it can comprehend adversarial and obfuscated samples. When more features are added, there is no interpretation’s fidelity, and also, there is no trust score to evaluate the extent of interpretability. LIME
Briguglio and Saad. 2019  [20] Aims to provide an interpretable malware detector to enhance the N-gram analysis They developed and evaluated the effectiveness of interpretability techniques using features of N-grams They interpreted the logistic regression model, single sample with random forest, neural network, single sample with neural network Systematic interpretation assessment for malware detection is not provided. Layer-wise relevance propagation (LRP)
Li et al. 2021  [65] Aims to provide excellent interpretability in malware classification through a machine learning model Proposed a novel machine learning model named LSH-base clustering approach to classifying the malware Their machine learning model generates assembly code function clusters, and the model itself supports the interpretation of the results Did not mention any drawbacks of the interpretable ML model Executable functions
Fidel et al. 2020  [37] Aims to detect evasion attacks such as adversarial examples by using the generated SHAP values as additional features The assumption of their work is that both the adversarial and benign samples will have different SHAP signatures A model learns and leverages the differences between the generated signatures and establishes the classifications It is not possible to transfer their attack on multiple detectors since the computation of SHAP values is not differentiable SHAP
Kumar and Subbiah 2022  [59] Aims to utilize static analysis for effective and explainable zero-day malware detection This study employs Shapley values to determine the most important features that are contributing more to predicting zero-day malware The SHAP explainability technique was utilized to determine the misclassification based on the topmost contributing features Reason for misclassification using SHAP should be explored more. SHAP
lee et al. 2022  [62] This work aims to provide a technique for screening high-quality data in order to detect false predictions based on reliability indicators. This study conducts experiments on two different datasets to analyze the effectiveness of their proposed approach The SHAP explainability technique was utilized to identify the most important features that are contributing to the classification so that even humans can easily interpret based on the AI predictions, and reliability is also calculated and analyzed Lacking the analysis on other explainability approaches such as LEMNA and LIME SHAP
Galli et al. 2024  [40] The paper aims to make AI-driven security tools more reliable and effective, especially in important security settings. The development and evaluation of an XAI framework specifically designed for behavioral malware detection. It assesses four XAI methods and three datasets to determine their effectiveness in making LSTM and GRU model decisions transparent. While the paper evaluates various XAI techniques on specific datasets, the generalizability of these methods to other types of malware or different datasets is not adequately addressed. SHAP, LIME, LRP, and attention mechanisms
@@
V Future Directions
Explainable machine learning is a rapidly evolving field with many ongoing challenges and opportunities for exploration. In the previous sections, we conducted an extensive review of various explainable machine-learning techniques, with a particular focus on malware classification and detection. however, as the application of explainable methods in malware detection becomes increasingly prevalent, new challenges continue to emerge. This section outlines several key challenges and potential research directions that researchers may pursue as future work in the area of explainable malware analysis and detection.
@@ V-A Improve Data Sets
Improving and updating malware datasets is a critical concern in the field of XAI. Many existing datasets are outdated and lack comprehensive coverage of current malware behaviors. These datasets often do not provide a sufficient volume of data for effectively training XAI applications. For instance, previous research on explainable Android malware detection primarily utilized the DREBIN dataset, which comprises 5,616 malicious instances and 121,329 benign instances [10]. This imbalance, where benign instances significantly outnumber malicious ones, can hinder the training of effective models.
Moreover, the size of the current datasets is generally too small to train robust models effectively. However, this field of research needs an unbiased, reasonably sized benchmark dataset that equally represents both benign and malicious behaviors. Accessing this kind of dataset is essential for evaluating explainable ML-based techniques and achieving reliable detection results. Furthermore, the DREBIN dataset, in particular, highlights the limitations of static analysis, pointing to the necessity for dynamic updates that support more comprehensive dynamic analyses. Additionally, there is potential for innovation in automated data generation and minimization techniques to accelerate the prediction process. For example, in hardware malware detection, researchers [105] have generated trace data to facilitate hardware trace analysis and distinguish between malware and benign programs. Future work could focus on enhancing these techniques to streamline and speed up the predictive capabilities of malware detection systems.
V-B Combining Static and Dynamic
@@
To address this gap, multiple ML classifiers will be leveraged to analyze both source code and runtime dynamic features. This dual approach aims to improve the efficiency and effectiveness of ML algorithms in distinguishing between benign and malicious applications. Moreover, to overcome the limitations in static and dynamic methodologies, researchers should also develop online and real-time explainable malware detection systems. These systems would continuously monitor the entire system to detect any possible malicious behavior or traces at any moment. Thus, the development of hybrid and online detection systems represents a significant research challenge in the field of explainable malware analysis.
Figure 14: Future Directions for Explainable Malware Detection.
@@ V-C Analyze Model-Agnostic Techniques
Explainable machine-learning techniques remain underutilized and underexplored in the domain of malware detection. There is a need to explore various model-agnostic techniques that provide both local and global explanations, which can help develop fast-training and explainable models without sacrificing accuracy. One promising direction for future research is the automation of explainability when it is decoupled from the underlying machine learning model. This decoupling facilitates the easy replacement of both the explainable technique and the machine learning model itself.
The model-agnostic technique LIME is widely cited in research but comes with notable limitations. Its reliance on data sampling can lead to variability in explanations, making them potentially unstable and unreliable. Furthermore, if the local fidelity measure is inaccurate, the reliability of LIME’s explanations for distinguishing between malicious and benign samples is compromised. Additionally, LIME lacks guidance on the optimal number of features to use, which could affect the quality of its explanations.
To advance the field of explainable malware detection, paying attention to evaluate various model-agnostic techniques is essential. Future research could also focus on improving fidelity in explanations, which is crucial for maintaining reliability in rapidly evolving scenarios. Overall, model-agnostic methods represent a flexible and effective approach to enhancing malware detection through explainable machine learning.
V-D Explainability Approach to Multiple Models
Previous research in explainable machine learning within the malware detection domain has primarily focused on developing frameworks and applying specific explainability methods to those frameworks. However, there is a notable gap in the literature regarding the selection of explainability techniques for non-differentiable models. Theoretical findings suggest that under certain assumptions, various machine learning algorithms can yield similar decision functions. This similarity raises a critical question: how does one select the most appropriate explainable technique for a given malware detection process?
@@ V-E Use Pre-Trained Models
Looking forward, in the malware detection domain, it would be advantageous to leverage existing explainable pre-trained models rather than building new models from scratch. This approach can save considerable time that would otherwise be spent collecting data and training models and enhances the efficiency of detecting malicious activities in systems that may already be compromised. In other words, utilizing pre-trained models can accelerate the deployment of malware detection systems and improve their effectiveness by integrating advanced, pre-learned features into the detection process. Applying this strategy could significantly extend the application of neural networks in malware detection which can lead to quicker adaptability and broader applicability.
@@ V-F Automated Explainable Models
One promising direction for future research in the field of malware detection involves implementing more automated explainability models. The goal is to enhance user trust in black-box models, such as those based on deep learning, which are currently not automatically interpretable. Most existing research focuses on interpreting the results of malware detection post hoc after the detection has already occurred. This method leaves a gap in real-time understanding and response, which automated explainability aims to fill.
Achieving an optimal balance between accuracy and explainability continues to be a significant challenge. Automated explainability could help bridge this gap by providing insights into the decision-making process of complex models in real-time. Additionally, there is a clear need for more research focused on quantitative-level evaluation of these explainable models. Such evaluations would assess the interpretability and how the introduction of explainability affects the overall performance of the detection system.
@@ V-G Improve detection efficiency
A valuable future direction in explainable malware detection is to enhance the design methodologies of malware detectors so that the explanations they generate can assist professionals in more accurately characterizing malware attacks, ultimately improving detection accuracy. For example, wang et al. [118] involve extracting features and employing a decision tree to develop a model capable of determining the maliciousness of applications.
@@
V-H Mitigating attacks
In recent research, the primary focus has been on black-box attacks, gradient-based attacks, evasion attacks, and poisoning attacks. Evasion attacks involve manipulating malicious input samples during the training phase to circumvent detection by a trained system, and it requires access to the model. Poisoning attacks compromise the integrity of training data by introducing incorrect data since it can mislead the learning process of ML models. This corruption of training data severely undermines the entire training process. In both gradient-based and poisoning attacks, it is assumed that the attacker has knowledge of the feature space used by the target. Future research in the field of explainable machine learning should explore defense mechanisms against these types of attacks and develop generic mitigation methods. Moreover, while current attacks typically use either static or dynamic approaches, future attacks might utilize a hybrid approach that integrates both strategies. As malware data continuously evolves, implementing attacks in online detection systems could pose significant challenges for attackers trying to intercept or manipulate high-speed continuous data compared to data stored on devices.

### v2

@@ IV-C Hardware-Based Malware Approaches
Many researchers have designed hardware-based malware detectors with the assumption that solutions such as anti-virus software can be fooled easily by malicious code. Although there are many studies on hardware-assisted malware detection, there is a distinct lack of research focused on explainability. This highlights a critical need for innovation in the field that goes beyond detection accuracy, aiming for systems that can articulate the rationale behind their detections. In table III, the literature on hardware malware classification and detection using XAI is discussed.
Sheldon [119] conducts a study analyzing hardware traces for malware detection using explainable ML approaches. Hardware traces comprise the data stored in caches and registers during the execution of a program. These traces are then processed by an ML-based model to identify malware presence. Subsequently, the model’s accuracy is evaluated against that of leading-edge ML models. The feedback obtained from this comparison is utilized to refine the accuracy of malware detection.
To explore the comprehensive study of hardware trace analysis and the development of an explainable, hardware-assisted malware detection framework, the research by Pan et al. [110] introduces a hardware framework depicted in Fig 5. This framework is structured around three principal activities.
@@
After determining the linear regression coefficients, the focus shifts to interpreting the outcomes, with particular emphasis on identifying the most influential features. Features associated with larger coefficients are flagged as potentially malicious. To benchmark the effectiveness of their method, the authors reference the PREEMPT malware detector for comparison [120]. PREEMPT employs two algorithmic models, i.e., Random Forest and Decision Tree, utilizing hardware performance counters (HPC) to generate its dataset. However, it does not extensively investigate the dataset’s characteristics, focusing primarily on the interpretation of outcomes. Given the scarcity of research in this area and the need to address malware detection challenges, it is imperative for researchers to engage deeply with the realm of explainable hardware malware detection.
Building upon their previous efforts, the authors [111] harness hardware performance counters and embedded trace buffers to identify the exact locations of malicious activities within a system. They develop Decision Tree and RNNs to perform a trade-off between accuracy and efficiency in their detection methodology. Their evaluation, conducted on a broad spectrum of real-world malware datasets, elucidates the interpretability of the RNN model, leveraging linear regression and Decision Tree through tree parsing techniques.
@@
In conclusion, hardware-based XAI solutions exploit HPCs and trace buffers to offer fine-grained inspection of runtime behavior. The interpretable models (e.g., linear regression, interpretable feed-forward nets) provide valuable insights into which low-level counters or basic blocks signal malicious patterns. However, custom hardware or specialized instrumentation is often needed, which can limit widespread deployment.
@@ IV-D PDF-Based Malware Approach
A Portable Document Format (PDF) file contains text, images, digital signatures, and other elements. Its structure includes a Header, Body, Cross-reference Table, and Trailer. The header of a PDF file is the top section that indicates the version number and file format. The body of the PDF stores all the pertinent data, and it contains a range of objects, including data, text, images, and dictionaries. The cross-reference table in a PDF file includes links to all elements within the document, which facilitates navigation and access. The trailer, which links to the cross-reference table, also contains the EOF marker.
@@
To classify PDF malware, another research [117] introduces an explainable method named LEMNA, which provides high-fidelity explanations for malware detection. They utilize deep learning models and assess their interpretability using LEMNA. The study also explores feature augmentation, along with synthetic and feature deduction tests. They note that due to sparse input feature vectors affecting local decision boundaries, LIME, and other advanced explainable techniques were as ineffective as traditional feature selection methods. The PDF-based malware detection studies are summarized in Table III.
PDF files continue to be a popular vector for malicious exploitation. The interpretability methods described here—whether gradient-based or model-agnostic—provide clarity on which structural elements (e.g., JavaScript objects, suspicious headers) are raising red flags. Nevertheless, sophisticated obfuscation within PDFs can still pose significant challenges for these models.
@@ IV-E Linux-Based Malware Approach
Wang et al. [113] focus on exposing vulnerabilities in malware detectors through explainability-guided evasion attacks that combine feature space manipulation with problem space obfuscation. They utilize a dataset of approximately 43,553 ELF binary files on Linux systems. Their research uses the model-agnostic explainability method SHAP to demonstrate how evasion attacks can be transferred from one detector to another. In another study, Mills et al. [114] develop a lightweight malware detection system named NODENS, suitable for deployment on Raspberry Pi hardware. They test several ML-based algorithms on a Linux operating system, with the Random Forest algorithm performing optimally among them. This work utilizes a tree-based model to facilitate visual interpretation of the classification process, which enhances the end user’s understanding of the output and aids in the individual development of the malware sample lifecycle. Due to the infrequent application of the Linux platform, there is limited research on explainable malware detection within the Linux domain. Studies addressing this topic are detailed in Table III.
These Linux-oriented approaches show that explaining which code snippets or system calls are pivotal can help analysts better grasp the root causes of malicious behavior. However, the limited availability of Linux malware datasets—and the complexity of dynamic sandboxing—remains a bottleneck for broad adoption.
@@ IV-F Other Approaches
TABLE IV: Research Addressing Explainable Machine Learning in Malware Analysis.
@@
Chen 2018 [122] Employ image-based static ML with LIME explanations Visualizes malicious features in grayscale images for improved interpretability Unverified against adversarial or obfuscated malware; lacks trust metrics LIME
Briguglio and Saad. 2019 [123] Use LRP to interpret ML-based n-gram analysis in malware detection Compared LR, RF, NN using LRP for n-gram feature relevance Lacks thorough interpretability framework analysis Layer-wise relevance propagation (LRP)
Li et al. 2021 [124] Propose LSH-based clustering for interpretable function-level malware classification Groups similar assembly functions, offering built-in interpretability Potential limitations not clearly discussed Executable functions
Fidel et al. 2020 [125] Exploit SHAP “signatures” to detect adversarial vs. benign samples Shows SHAP-value distributions differ for adversarial samples, enabling classifier-based detection Method not easily transferable across detectors; SHAP’s non-differentiability is limiting SHAP
Kumar and Subbiah 2022 [126] Apply SHAP to identify top features in zero-day malware classification Employs SHAP bar/waterfall plots to analyze false positives/negatives in zero-day detection Does not deeply explore SHAP-based misclassification causes SHAP
lee et al. 2022 [127] Use reliability indicators + SHAP to refine alerts in large-scale security threats Validated screening method with IDS/malware datasets, employing SHAP-driven feature insights Omits comparison with alternate XAI tools SHAP
Galli et al. 2024 [128] Evaluate SHAP, LIME, LRP, attention to explain LSTM/GRU in behavioral malware detection Proposed an XAI framework for behavioral models, assessing multiple explanation methods across diverse datasets Unclear if results generalize beyond tested datasets SHAP, LIME, LRP, and attention mechanisms
@@
These generalized XAI solutions—covering n-gram analysis, signature-based detection, or adversarial defense—offer additional perspectives on explaining ML-driven malware detection. While each technique addresses a different niche (e.g., zero-day detection, adversarial resilience, multi-platform coverage), common XAI issues like scalability, consistency of explanations, and susceptibility to manipulation still arise.
In summary, explainable malware detection strategies span a diverse range of platforms and techniques—from gradient-based methods pinpointing critical file regions to model-agnostic tools offering broader coverage at the cost of local-only explanations to advanced image-based or rule-based systems. While these approaches strengthen trust, transparency, and analyst insight, they also face challenges in scaling to large datasets, handling obfuscated or adversarial samples, and balancing interpretability with accuracy. In the next section, we examine how these XAI methods can be enhanced against advanced threat tactics and integrated into practical cybersecurity frameworks.
V Future Research Directions
Explainable ML is an evolving field with many ongoing challenges and opportunities for exploration. In the previous sections, we conducted an extensive review of various explainable ML techniques, with a particular focus on malware classification and detection. however, as the application of explainable methods in malware detection becomes increasingly prevalent, new challenges continue to emerge. As shown in Fig 6, this section outlines several key challenges and potential research directions that researchers may pursue as future work in the area of explainable malware analysis.
Figure 6: Future Directions for Explainable Malware Detection.
@@ V-A Improve Datasets
Improving and updating malware datasets is a critical concern in the field of XAI. Many existing datasets are outdated and lack comprehensive coverage of current malware behaviors. These datasets often do not provide a sufficient volume of data for training XAI applications. For instance, previous research on explainable Android malware detection utilized the DREBIN dataset, which comprises 5,616 malicious instances and 121,329 benign instances [107]. This imbalance, where benign instances significantly outnumber malicious ones, can hinder the training of effective models.
Moreover, the size of the current datasets is generally too small to train robust models. This field of research needs an unbiased, reasonably sized benchmark dataset that equally represents both benign and malicious behaviors. Accessing this kind of dataset is essential for evaluating explainable ML-based techniques and achieving reliable detection results. Furthermore, the DREBIN dataset, in particular, highlights the limitations of static analysis, pointing to the necessity for dynamic updates that support more comprehensive dynamic analyses. Additionally, there is potential for innovation in automated data generation and minimization techniques to accelerate the prediction process. For example, in hardware malware detection, researchers [119] have generated trace data to facilitate hardware trace analysis and distinguish between malware and benign programs. Future work could focus on enhancing these techniques to streamline and speed up the predictive capabilities of malware detection systems.
V-B Combine Static and Dynamic
@@
To address this gap, multiple ML-based classifiers will be leveraged to analyze both source code and runtime dynamic features. This dual approach aims to improve the efficiency and effectiveness of ML-based algorithms in distinguishing between benign and malicious applications. Moreover, to overcome the limitations in static and dynamic methodologies, researchers should also develop online and real-time explainable malware detection systems. These systems would continuously monitor the entire system to detect any possible malicious behavior or traces at any moment. Thus, the development of hybrid and online detection systems represents a significant research challenge in the field of explainable malware analysis.
@@ V-C Analyze Model-Agnostic Techniques
There is a need to explore various model-agnostic techniques that provide both local and global explanations, which can help develop fast-training and explainable models without sacrificing accuracy. One promising direction for future research is the automation of explainability when it is decoupled from the underlying machine learning model. This decoupling facilitates the easy replacement of both the explainable technique and the machine learning model itself.
The model-agnostic technique, LIME, is widely cited in research but comes with notable limitations. Its reliance on data sampling can lead to variability in explanations, making them potentially unstable and unreliable. Furthermore, if the local fidelity measure is inaccurate, the reliability of LIME’s explanations for distinguishing between malicious and benign samples is compromised. Additionally, LIME lacks guidance on the optimal number of features to use, which could affect the quality of its explanations.
To advance the field of explainable malware detection, paying attention to evaluating various model-agnostic techniques is essential. Future research could also focus on improving fidelity in explanations, which is crucial for maintaining reliability in rapidly evolving scenarios. Overall, model-agnostic methods represent a flexible and effective approach to enhancing malware detection through explainable ML.
V-D Apply Explainability Approach to Multiple Models
Previous research in explainable ML within the malware detection domain has primarily focused on developing frameworks and applying specific explainability methods to those frameworks. However, there is a notable gap in the literature regarding the selection of explainability techniques for non-differentiable models. Theoretical findings suggest that under certain assumptions, various ML-based algorithms can yield similar decision functions. This similarity raises a critical question: how does one select the most appropriate explainable technique for a given malware detection process?
@@ V-E Use Pre-Trained Models
Looking forward, in the malware detection domain, it would be advantageous to leverage existing explainable pre-trained models rather than building new models from scratch. This approach can save considerable time that would otherwise be spent collecting data and training models and enhances the efficiency of detecting malicious activities in systems that may already be compromised.
In other words, utilizing pre-trained models can accelerate the deployment of malware detection systems and improve their effectiveness by integrating advanced, pre-learned features into the detection process. At present, there is no well-established, publicly available pre-trained explainable model that is widely adopted specifically for malware detection. While the concept of pre-trained” and “explainable models exists in other domains (e.g., NLP, computer vision), the malware analysis community has not converged on a standard, large-scale pre-trained model that includes built-in explainability for detecting malicious code or behavior.
Building a comprehensive pre-trained explainable model for malware detection is hampered by challenges such as limited access to large, high-quality datasets due to privacy and proprietary concerns, as well as the constantly evolving nature of malware which requires frequent retraining and adaptation. Additionally, integrating explainability mechanisms at scale introduces complexity and may reduce performance if not carefully designed and maintained.
@@ V-F Automated Explainable Models
One promising direction for future research in the field of malware detection involves implementing more automated explainability models. The goal is to enhance user trust in black-box models, such as those based on DL techniques, which are currently not automatically interpretable. Most existing research focuses on interpreting the results of malware detection after the detection has already occurred. This method leaves a gap in real-time understanding and response, which automated explainability aims to fill.
Moreover, achieving an optimal balance between accuracy and explainability continues to be a significant challenge. Automated explainability could help bridge this gap by providing insights into the decision-making process of complex models in real time. Additionally, there is a clear need for more research focused on quantitative-level evaluation of these explainable models. Such evaluations would assess the interpretability and how the introduction of explainability affects the overall performance of the detection system.
@@ V-G Improve Detection Efficiency
A valuable future direction in explainable malware detection is to enhance the design methodologies of malware detectors so that the explanations they generate can assist professionals in more accurately characterizing malware attacks. For example, wang et al. [93] involve extracting features and employing a decision tree to develop a model capable of determining the maliciousness of applications.
@@
V-H Mitigate attacks
In recent research, the primary focus has been on black-box attacks, gradient-based attacks, evasion attacks, and poisoning attacks. Evasion attacks involve manipulating malicious input samples during the training phase to circumvent detection by a trained system, and it requires access to the model. Poisoning attacks compromise the integrity of training data by introducing incorrect data since it can mislead the learning process of ML models. This corruption of training data severely undermines the entire training process.
In both gradient-based and poisoning attacks, it is assumed that the attacker has knowledge of the feature space used by the target. Future research in the field of explainable ML should explore defense mechanisms against these types of attacks and develop generic mitigation methods. Moreover, while current attacks typically use either static or dynamic approaches, future attacks might utilize a hybrid approach that integrates both strategies. As malware data continuously evolves, implementing attacks in online detection systems could pose significant challenges for attackers trying to intercept or manipulate high-speed continuous data compared to data stored on devices.